### v3

# Is Hyper-Parameter Optimization Different for Software Analytics?

Abstract
Yes. SE data can have “smoother” boundaries between classes (compared to traditional AI data sets). To be more precise, the magnitude of the second derivative of the loss function found in SE data is typically much smaller. A new hyper-parameter optimizer, called SMOOTHIE, can exploit this idiosyncrasy of SE data. We compare SMOOTHIE and a state-of-the-art AI hyper-parameter optimizer on three tasks: (a) GitHub issue lifetime prediction (b) detecting static code warnings false alarm; (c) defect prediction. For completeness, we also show experiments on some standard AI datasets. SMOOTHIE runs faster and predicts better on the SE data–but ties on non-SE data with the AI tool. Hence we conclude that SE data can be different to other kinds of data; and those differences mean that we should use different kinds of algorithms for our data. To support open science and other researchers working in this area, all our scripts and datasets are available on-line at https://github.com/yrahul3910/smoothness-hpo/.
Index Terms:
software analytics, smoothness, hyper-parameter optimization
1 Introduction
To say the least, much current work in SE uses algorithms from the AI community. Usually, these algorithms are used in their default “off-the-shelf” configuration [23, 67].
Is this unwise? Is software engineering data fundamentally the same as those used to develop those AI tools? Or does software engineering datasets have unique characteristics that makes it different from others, and therefore would benefit from other kinds of AI tools?
This article argues that it is not always advisable to apply standard AI tools to SE data. Specifically, in the case of hyper-parameter optimization (HPO), we show that:
  • AI for SE is different to standard AI. SE data can have much “smoother” boundaries between classes. To be precise, compared to AI data, the magnitude of the second derivative of the loss function found in SE data can be much smaller.
  • For SE data, better HPO result can be achieved using tools designed for SE data.
This is an important result since HPO is an important problem. Hyper-parameter optimizers explore the myriad of configuration choices which control either the learning or the data pre-processing. As shown by Table I and Table II, there are many such hyper-parameters. Much SE research has shown that HPO is very effective at improving the performance of models built from SE data [24, 2, 70, 72]; for example, see Figure 1. Nevertheless, while studying industrial text-mining tools, we have found many examples of poorly-chosen hyper-parameters (which meant our industrial partners were selling less-than-optimal products) [41].
Figure 1: Defect prediction: likelihood of a learner performing best. BLUE shows performance pre-optimization and RED shows very large performance improvements (post-optimization). E.g. on the left-hand-side, the C5.0 decision tree learner had worst performance (pre-optimization) and best performance (post-optimization). From [58].
This paper offers a new hyper-parameter optimizer called SMOOTHIE learned from SE data that selects hyper-parameter options according to how well those options simplify the boundary between (say) defective and non-defective modules. For this paper, the important features of SMOOTHIE are that:
  • For SE data, SMOOTHIE makes better predictions than state-of-the-art-methods taken from AI;
  • SMOOTHIE runs much faster than that state-of-the-art;
  • But SMOOTHIE performs poorly for non-SE data.
To say all that another way:
SE data can be different to other kinds of data; and those differences mean that we should use different kinds of algorithms for our data.
The rest of this paper is structured as follows. The next section discusses how SE data is different to AI data. This is followed by notes on related work, specifically, hyperparameter optimization and its role in software analytics. Section 5 discusses our experimental setup, and the results of those experiments are presented in Section 6. We discuss why SMOOTHIE works so well for SE data along with threats to validity in Section 7, before discussing limitations and concluding.
2 Unique Properties of SE Data
Our premise is that SE needs its own kind of AI since data from SE projects is somehow fundamentally different (than the data used to develop standard AI algorithms). This section offers evidence for that premise, in two parts. Firstly we will argue that we expect SE data to be smooth. Secondly, we will also say that other kinds of data may not be as smooth.
2.1 Expect Smoothness in SE Data
A repeated result in SE is that SE artifacts are surprisingly regular. Hindle et al. [31] writes that “programming languages, in theory, are complex, flexible and powerful, but the programs that real people write are … rather repetitive, and thus they have statistical properties that can be captured”.
There are many possible reasons for this repetition. For example, successful software is maintainable; i.e. it contains structures recognizable to those doing the maintenance. Hence there are selection pressures forcing those who write software to make repeated use of:
  • Readily identified idioms such as for-loops;
  • Or higher level patterns such as subject-observer [25];
  • Or even higher-level architectural patterns such as the microservices architecture discussed by [14].
Such repetition means that the state space of software can contain many repeated structures– which would mean that different parts of a code base may be surprisingly similar. Data generated from such repetitive structures might itself be repetitive (with minimal differences between adjacent clusters in the data).
Whatever the reason for this repetition, that regularity would mean there are few differences between adjacent parts of the code. Such similarities would manifest as smooth surfaces within the data. Hence, for some time now, the authors have been asking if we can see such smoothness and can we use that smoothness to achieve some useful results.
It was therefore gratifying, in Fall 2023, to actually observe such smoothness. While working on static code analysis false alarm identification and hyper-parameter optimization [72], one of us (Yedida) noticed a pattern in changes to the loss functions before and after hyper-parameter optimization. Loss functions are surrogates for metrics that have gradients (and the better the learner, the lower the loss, generally speaking). Figure 2 left and right shows that:
After optimization, the loss function gets “smoother” (i.e. on average, fewer changes in any local region).
To say that another way, a learner that respects the smoothness of SE data does better than otherwise. Based on that observation, we theorized that:
Hyper-parameter optimization is best viewed as a “smoothing” function for the decision landscape.
Figure 2: Left & right shows loss landscape before & after HPO. The loss function be visualized as the $z$ dimension of domain data mapped into two PCA components $x,y$. In this figure, red indicates regions with highest cross-entropy error.
This theory predicts that the greater the smoothing, the better the hyper-parameter optimization. This paper tests that prediction. The algorithm we call SMOOTHIE is an experiment on whether or not increasing smoothness could be used to guide hyper-parameter optimization (HPO). As shown in this paper, SMOOTHIE can assess hyper-parameter options without needing to run a learner to completion. For example, for neural networks that learn over 100 epochs, SMOOTHIE can assess hyper-parameter options within the first epoch. That is, SMOOTHIE can focus on good configurations by very quickly rejecting bad configurations using a fraction of the CPU cost of conventional hyper-parameter optimizers.
2.2 Other Kinds of Data May be Less Smooth
We argued above that SE practices are designed to generate maintainable software, and that such practices may also generate SE data that exhibits smoothness. Data from some non-SE sources is not constrained in the same way. Hence, we would expect non-SE data to be less smooth than SE data.
This is actually the case. Figure 3 shows the smoothness measured in the data used in this study. This figure was calculated using some equations described later in this paper. In summary, smoothness is a measure of the magnitude of the second derivative in the loss function. Lower smoothness means more regularity and repeated properties between those adjacent parts [^1] Although unintuitive, the mathematical term “smoothness” is a bound on the sharpness as opposed to the intuitive smoothness. We will make this precise later..
In Figure 3, blue shows the smoothness of the loss function for the SE data mentioned in the last section. Also in that figure, orange shows the smoothness seen in various AI data sets from UCI (the data sets here were selected from prior work that also explored the value of different algorithms for SE and non-SE dats [3]).
The stand-out feature of this graph is that the SE data has dramatically different smoothness to the non-SE data. This paper will show that the x-axis of this figure can predict what algorithm will be most successful:
  • At high smoothness values, SMOOTHIE loses to standard AI algorithms.
  • But, as we shall show, SMOOTHIE defeats them at low smoothness values.
Clearly, this result supports the claim in the introduction that (a) SE data can be different to standard AI data so (b) it is unwise to always assume that AI algorithms should be applied to SE data.
3 Hyper-parameter Optimization (HPO)
We said above that low/high smoothness can predict for our new algorithm winning/losing against standard AI algorithms. The task in that experimental comparison was hyperparameter optimization (HPO). This section defines HPO, and offers examples of how it is used in SE. Afterwards, the rest of this paper uses HPO to show the value of smoothness.
Figure 3: Distribution of smoothness for SE and non-SE datasets. Calculated using the maths of §3. See §3.2 for the data used in this figure. For details on how the x-axis values are calculated, see §4. Note the mean values for the two populations are different: mean smoothness = 2, 5 for SE, AI respectively.
TABLE I: A sample hyper-parameter choices for learners (note, this is not a complete set).
Feedforward networks might have (say) 3 to 20 layers and 2 to 6 units per layer
Deep learners have many more control parameters including how to set their topology and activation functions.
Neighbor classifiers need to know (a) how to measure distance; (b) how many neighbors to use; (c) how to combine those neighbors; (d) how (or if) to pre-cluster the data (to optimizer th neighborhood lookup); etc.
Random forests need to know how many trees to build, how many features to use in each tree, and how to poll the whole forest (e.g., majority or weighted majority).
Decisions for logistic regression include which regularization terms ($L_{1}$, $L_{2}$) to use, (or, indeed, to use both) and (b) the strength of the regularization; i.e. $C\in\{0.1,1,10\}$.
TABLE II: A sample of hyper-parameter choices for data pre-processing (note, this is not a complete set).
Some of pre-processors are simple: normalization: $x\to\frac{x}{\lVert x\rVert}$; standardization: $x\to\frac{x-\mu_{x}}{\sigma_{x}}$; min-max scaling: $x\to\frac{x-\min x}{\max x-\min x}$; max-absolute value scaling: $x\to\frac{x}{\max\lvert x\rvert}$); robust scaling: $x\to\frac{x-P_{50}(x)}{P_{75}(x)-P_{25}(x)}$ where $P_{i}(x)$ represents the $i$th percentile of x. Various authors recommend one or more of these simpler pre-processors. For example, for instance-based methods Aha et al. [4] recommend min-max scaling since it means numbers on different scales can be compared.
Another widely used pre-processor is SMOTE [13]. In order to handle unbalanced training data, SMOTE generates synthetic minority samples by extrapolating along line segments joining the nearest neighbors of the minority class.
“Fuzzy sampling” [70], concentrically adds $\frac{1/n}{2^{i}}$ samples at concentric layer $i$, for $\left\lfloor{\log_{2}{1/n}}\right\rfloor$ layers, where $n$ is the class imbalance ratio. Unlike other oversampling techniques, fuzzy sampling reverses the class imbalance. The authors show that applying it twice can also be beneficial.
Another pre-processor from semi-supervised learning [72] (which we call “label engineering”) is to delete, at random $m-\sqrt{m}$ labels. These are then recovered by building a kd-tree [6] on the remaining $\sqrt{m}$ samples. For each lost label, the kd-tree is queried for the nearest $4]{m}$ samples, and the mode of their labels is assigned.
3.1 About HPO
A learning system combines some data pre-processing and machine learning. Every such system has magic control parameters that adjust the data pre-processing and/or the machine learning (e.g. see Table I and Table II). Even something as simple as the $k-$nearest neighbors classifier needs to decide how many neighbors it should study in order to classify a new example (and, as seen in Table I, it must settle on several other configuration options).
In essence, HPO is the process of fiddling with the control parameters of a learner and the data pre-processing. The trivial way to do this is grid search, which is a set of nested for-loops across all possible values of all the control parameters. In practice, grid search can be impractically slow and ineffective [7].
Figure 4: A brief overview of hyper-parameter optimization methods. For a more in-depth review, see [57, 22, 9].
Smarter HPO algorithms take more care with how they explore the hyper-parameter space. For example, [8, 7] discuss methods to reduce the cost of hyper-parameter optimization via some clever random sampling. They demonstrate the effectiveness of random search under a limited computational budget, additionally showing that for most problems, only a few hyper-parameters matter, although that subset of important hyper-parameters is different across datasets. To introduce the prior state-of-the-art, we now discuss two more strategies to explore the hyper-parameter space more efficiently.
Bayesian optimization [52] is a sequential model-based optimization algorithm that improves as it gathers more data. It involves a surrogate modeling for the objective function and an acquisition function that samples configurations. Surrogates are a model learnt from just a few examples which are used to assess all the examples. Surrogates serve as a cheap “acquisition functions” to find interesting examples of what to evaluate next.
A widely-used surrogate method is Gaussian Process (GP) models [64]. Various acquisition functions have been proposed, such as Expected Improvement [36], Gaussian Process Upper Confidence Bound (GP-UCB) [42], and predictive entropy search [30]. Among these, Expected Improvement is the most popular choice. Of the various BO-based methods, BOHB [19] combines Bayesian optimization and Hyperband [^2] Briefly, Hyperband is based on “successive halving”, a method that eagerly discards unpromising candidates at different computational budgets (so-called “fidelities”)–hence the term “multi-fidelity” is used to describe it. [44], and has emerged as the most popular surrogate-based algorithm combining these two due to it being open-source and state-of-the-art. The BO closely resembles TPE [8] with a single multi-dimensional kernel density estimate [73]. In a large-scale study, Eggensperger et al. [18] endorsed multi-fidelity methods (such as BOHB) over their black-box counterparts in constrained budget settings. This result was corroborated by a more recent work by Pfisterer et al. [55], which showed that in constrained budget settings, multi-fidelity methods outperformed full-fidelity methods such as SMAC [46].
Another strategy for searching in a more informed manner is differential evolution (DE). This is an evolutionary method that maintains a population of candidate solutions at each generation, and uses three randomly chosen individuals to build a new candidate for the next generation. In fact, DE has already seen applications in the SE literature: Majumder et al. [48] showed that using DE to tune fast classical learners proved to be over 500 times faster than deep learning, while outperforming it; and Agrawal and Menzies [1] showed that using DE to tune SMOTE [13] (an oversampling algorithm) can significantly improve the performance of many learning algorithms. DEHB [5] uses DE on top of Hyperband. On top of improving its performance, it allows for a distributed tuning of hyper-parameters.
HPO is a very large area of research and this section has only sampled some of that work. Figure 4 shows a broad classification of HPO algorithms, along with a few notable examples. For further details on HPO, see [57, 22, 9].
Many studies from the AI literature endorse BOHB and DEHB as the current state-of-the-art [73, 68, 22]. Hence, when we need to compare SMOOTHIE against some state-of-the-art AI HPO algorithm, we will use these two, along with random search as a simpler baseline.
3.2 HPO in software analytics
Here, we offer details of some case studies with SE and HPO. The following list is hardly exhaustive, but it does detail some of the more recent work in this area. Also, this section serves to introduce the examples used later in this paper.
3.2.1 Defect prediction.
Software quality assurance budgets are finite while assessment effectiveness increases exponentially with assessment effort [24]. Therefore, standard practice is to apply slower methods on code sections that seem most interesting (most critical or bug-prone). Defect prediction is a fast way to learn these interesting regions using static code attributes. A 2018 study [61] of 395 practitioners across 33 countries found that over 90% were willing to adopt defect prediction models in their workflow.
As shown in Figure 1, many of the algorithms used in defect prediction can benefit from hyper-parameter optimization. Recently, Yedida and Menzies [70] proposed “fuzzy sampling” (see Table II) and combined it with stacked autoencoders, SMOTE, hyper-parameter optimization (using DODGE [2]), and a weighted loss function to achieve state-of-the-art results on defect prediction using the PROMISE repository [10], which contains 20 static code attributes of the software such as the depth of inheritance tree [15], and mean McCabe cyclomatic complexity [51]. In this study, we use the same 6 datasets as that study for a fair comparison.
3.2.2 Issue lifetime prediction.
Issue lifetime predictors learn predictors for how long it takes to address issues raised by other developers (or end-users). Predicting issue close time has multiple benefits for developers, managers, and stakeholders, including task prioritization and resource allocation. The task is interesting since in open source software, developers triage their work according to the difficulty of their work; i.e. they will explore many small tasks before tackling harder tasks that are slower to accomplish.
Hyper-parameter optimization has been successfully applied to issue lifetime prediction. For example, recently we studied the DeepTriage [49] system. This is a state-of-the-art-tool that uses a bidirectional LSTM with attention after removing special characters, tokenization, and pruning. That code base is complex and took eight hours to train a predictor. Recently we found [71] that DeepTriage can be defeated using HPO and a forty-year old feed-forward neural net. The reason for this superior performance was simple: DeepTriage and that older neural net need eight hours and three seconds (respectively) to build a predictor. The faster algorithm is hence (much) more suitable for multiple hyper-parameter optimization studies. Hence the conclusion of that study was that "old, fast, and simpler algorithms (plus HPO) can defeat new, slow, and supposedly smarter algorithms”.
This paper uses the same data sets as that DeepTriage study; i.e. data mined from Bugzilla for three popular software projects: Firefox, Chromium, and Eclipse. For simplicity, although the data contains user comments and system records (both of which are text data sources), we do not use those features in this study for our method. Lee et al. [43] proposed binning the data into $k\in\{2,3,5,7,9\}$ classes and test the performance of classifiers under the binary and multi-class settings. For brevity, we show results for 2 and 3 classes.
3.2.3 Static code analysis false alarms
Static analysis (SA) tools report errors in source code, without needing to execute that code (this makes them very popular in industry). SA tools generate many false alarms that are ignored by practitioners. Previous research work shows that 35% to 91% of SA warnings reported as bugs by SA tools are routinely ignored by developers [29, 28]. Hence many researchers add a machine learning post-processor that reports the “alarms you should not ignore” (learned from the historical log of what errors do/do not result in developers changing the code) [62, 69, 72].
Initial results from Yang et al. [69] suggested that it was possible to find the actionable static code warnings generated by a static analysis tool like FindBugs, using a very simple learner. But later work by Kang et al. [38] showed that the datasets used by the former study had errors, and once manually repaired, those simple methods did not produce good results. This was later fixed by Yedida et al. [72], using hyper-parameter optimization that explored combinations of pre-processing methods. While successful, the methods of Yedida et al. were somewhat of a hack. The maths and methods of this paper replaces that hack with an approach that is much more well-founded approach (and which is simpler to apply).
4 About SMOOTHIE
The last section described some of the case studies we will explore (defect prediction, issue lifetime prediction, and static code analysis false alarms). The experiments of this paper check if there is any value in turning “bumpy” landscapes (like Figure 2, left-hand-side) into “smoother” landscapes (like Figure 2, right-hand-side). To conduct such experiments, we used the “smoothing” tool described in this section.
SMOOTHIE begins by sampling a few configurations of a learner, selected at random from, say, Table II. Then, we compute the smoothness for each configuration using the methods described in this section.
Input : Number of configs to sample $N_{1}$, number of configs to run $N_{2}$. Defaults: $N_{1}=30,N_{2}=5$.
Output : A near-optimal configuration
1 $\mathcal{H}_{0}\leftarrow$ Random($\mathcal{H},N_{1}$);
2 $S\leftarrow\phi$;
3 $P\leftarrow\phi$;
4 for each config $h$ in $\mathcal{H}_{0}$ do
5       S[config] $\leftarrow$ Get-Smoothness(h)
6 end for
7for each config in Top(S, $N_{2}$) do
8       P[config] $\leftarrow$ Run(config);
9
10 end for
11
12return $\operatorname*{arg\,max}P$;
Algorithm 1 SMOOTHIE
Algorithm 1 shows the pseudo-code for SMOOTHIE. We first pick $N_{1}$ ($=30$) configurations at random. For each of those configurations, we compute the smoothness and store it. Next,we pick the top $N_{2}$ ($=5$) of those configurations by smoothness values, and run only those, getting the performance for each. The algorithm then return the best-performing configuration.
The rest of this section shows the Get-Smoothness calculation used in Algorithm 1. But before that, we make the following important points. Our definition of smoothness is general to many learners (See Equation 3). However, that definition must be realized in different ways for different learners.
  • In some cases smoothness can be computed just by looking at certain properties of the data (in Gaussian Naive Bayes, that is the covariance matrix of the data, see Theorem 2). In this first case, HPO is very fast–just peek at the data to rank and prune the operators of (say) Table II.
  • In other cases, as in neural networks (see Theorem 9), smoothness changes by the inference process. In this latter case, we have to watch the learner running (at least for a limited time) before smoothness can be computed [^3] This brings the weights closer to the final weights at an exponential rate, assuming the loss is also strongly convex..
The following assumes some familiarity with the math behind neural networks, and some comfort with linear algebra and multivariate calculus. Readers who are interested in the results can, with no loss of continuity, merely read the statements of the theorems (skipping the lemmas) and the last paragraphs of each subsection to understand how the smoothness is computed. Our code is also open-sourced, and some readers might find it easier to instead read the implementation.
4.1 Preliminaries
Our description of SMOOTHIE uses the following notation.
For any learner, $X,y$ will represent a training set containing multiple independent values $X$ and a single goal $y$. In neural networks, we also use $W$ to denote the weights on the links between neurons. We also say:
  • $m$ is the number of training samples;
  • $n$ is the number of features;
  • $k$ is the number of classes, i.e. possible values for $y$;
  • $E$ denotes the loss function.
For a feedforward neural network, $L$ represents the number of layers. In that learner, at each layer, the following computation is performed:
$z^{[l]}$ $=W^{[l]T}a^{[l-1]}+b^{[l]}$ (1)
$a^{[l]}$ $=g^{[l]}(z^{[l]})$
$a^{[0]}$ $=X$
Here, $b^{[l]}$ is the bias vector at layer $l$, $W^{[l]}$ is the weight matrix at layer $l$, and $g^{[l]}$ is the activation function used at layer $l$ (we assume ReLU $g(x)=\max(0,x)$). The result from applying the activation function is called the activations at layer $l$, and is denoted by $a^{[l]}$. In this text, $\nabla_{W}f$ refers to the first gradient of $f$, with respect to $W$. Finally, $\nabla^{2}_{W}f$ denotes the second gradient, also called the Hessian. We use $\sup f$ to denote the least upper bound (supremum) of $f$.
An interesting use case for the gradient is the $L-$Lipschitzness, which can be defined as $\sup\lVert\nabla_{W}f\rVert$ [^4] In fact, it can be shown that a differentiable function $f$ is $L-$Lipschitz iff $\lVert\nabla f\rVert\leq L$ More strictly, a function $f$ is said to be $L-$Lipschitz if there exists a constant $L$ such that
$\centering\frac{\lVert f(y)-f(x)\rVert}{\lVert y-x\rVert}\leq L\@add@centering$ (2)
for all $x,y$ in the domain of $f$. The Lipschitz constant gives an idea of the curvature bound of the function under review.
The second gradient (the Hessian) is denoted by $\mathcal{H}=\nabla^{2}f$. In the strictest sense, the Hessian is a square tensor of second-order partial derivatives, and is symmetric (by Schwarz’s theorem). A continuously differentiable function $f$ is said to be $\beta-$smooth if its gradient is $\beta-$Lipschitz, or
$\frac{\lVert\nabla f(y)-\nabla f(x)\rVert}{\lVert y-x\rVert}\leq\beta$ (3)
From the mean-value theorem, there exists $v$ such that
$\frac{\lVert\nabla f(y)-\nabla f(x)\rVert}{\lVert y-x\rVert}=\lVert\nabla^{2}f(v)\rVert\leq\sup\limits_{v}\lVert\nabla^{2}f(v)\rVert$
(Note that this computation assumes the loss function is twice-continuously differentiable.)
Hence, $\sup\limits_{v}\lVert\nabla^{2}f(v)\rVert$ is such a $\beta$ [^5] Since all norms are equivalent in finite dimension, the choice of the norm (such as Euclidean or Frobenius) of the Hessian is irrelevant for our application, and different configurations can be ordered by smoothness as long as the choice of norm remains consistent.. Since $\beta$ is the least such constant, $\beta\leq\sup\limits_{v}\lVert\nabla^{2}f(v)\rVert$.
4.2 Smoothness for Feedforward Classifiers
This section shows the smoothness computation for different learners. We prove the result for feedforward classifiers here; for other derivations (Naive Bayes, logistic regression), please see the Supplementary Material.
First, we will prove an auxiliary result that will support our main proof. Readers who are interested in the equation for the smoothness can skip ahead to Theorem 9.
Lemma 1.
For a deep learner with ReLU activations in the hidden layers, the last two layers being fully-connected, and a softmax activation at the last layer,
$\dfrac{\partial E}{\partial z_{j}^{[L]}}\leq\frac{k-1}{km}\left(\sum\limits_{i=1}^{m}[y^{(i)}=j]\right)$
under the cross-entropy loss.
Proof.
We will use the chain rule, as follows:
$\frac{\partial E}{\partial z^{[L]}_{j}}$ $=\frac{\partial E}{\partial a^{[L]}_{j}}\cdot\frac{\partial a^{[L]}_{j}}{\partial z^{[L]}_{j}}$ (4)
Consider the Iverson notation version of the general cross-entropy loss:
$E(\textbf{a}^{[L]})=-\frac{1}{m}\sum\limits_{i=1}^{m}\sum\limits_{h=1}^{k}[y^{(i)}=h]\log a_{h}^{[L]}$
Then the first part of (4) is trivial to compute:
$\frac{\partial E}{\partial a_{j}^{[L]}}=-\frac{1}{m}\sum\limits_{i=1}^{m}\frac{[y^{(i)}=j]}{a_{j}^{[L]}}$ (5)
The second part is computed as follows.
$\frac{\partial a^{[L]}_{j}}{\partial z^{[L]}_{p}}$ $=\frac{\partial}{\partial z^{[L]}_{p}}\left(\frac{e^{z^{[L]}_{j}}}{\sum_{l=1}^{k}e^{z^{[L]}_{l}}}\right)$
$=\frac{[p=j]e^{z^{[L]}_{j}}\sum_{l=1}^{k}e^{z^{[L]}_{l}}-e^{z^{[L]}_{j}}\cdot e^{z^{[L]}_{p}}}{\left(\sum_{l=1}^{k}e^{z^{[L]}_{l}}\right)^{2}}$
$=\frac{[p=j]e^{z^{[L]}_{j}}}{\sum_{l=1}^{k}e^{z^{[L]}_{l}}}-\frac{e^{z^{[L]}_{j}}}{\sum_{l=1}^{k}e^{z^{[L]}_{l}}}\cdot\frac{e^{z^{[L]}_{p}}}{\sum_{l=1}^{k}e^{z^{[L]}_{l}}}$
$=\left([p=j]a^{[L]}_{j}-a^{[L]}_{j}a^{[L]}_{p}\right)$
$=a^{[L]}_{j}([p=j]-a^{[L]}_{p})$ (6)
Combining (5) and (6) in (4) gives
$\frac{\partial E}{\partial z^{[L]}_{j}}=\frac{1}{m}\left(\sum\limits_{i=1}^{m}[y^{(i)}=j]\right)\left(\sum\limits_{h=1}^{k}a^{[L]}_{j}-[h=j]\right)$ (7)
As a prudence check, note that $E$ is a scalar and $z^{[L]}_{j}$ is a vector (over all the samples); the right-hand side is also a vector (by vectorizing the Iverson term $[h=j]$). It is easy to show that the limiting case of this is when all softmax values are equal and each $y^{(i)}=p$; using this and $a^{[L]}_{j}=\frac{1}{k}$ in (7) and combining with (4) gives us our desired result:
$\sup\limits_{j}\left\lVert\frac{\partial E}{\partial z^{[L]}_{j}}\right\rVert=\frac{k-1}{km}\left(\sum\limits_{i=1}^{m}[y^{(i)}=j]\right)$ (8)
Theorem 1 (Smoothness for feedforward networks).
For a deep learner with ReLU activations in the hidden layers, the last two layers being fully-connected, and a softmax activation at the last layer, the smoothness of the cross-entropy loss is given by
$\sup\lVert\nabla^{2}_{W}E\rVert\propto\frac{k-1}{km}\sup\frac{\lVert a^{[L-1]}\rVert}{\lVert W^{[L]}\rVert}$ (9)
Proof.
Lemma 1 gives us (using the chain rule one step further):
$\nabla_{W}E\leq\frac{(k-1)}{km}\left(\sum\limits_{i=1}^{m}[y^{(i)}=j]\right)a_{j}^{[L-1]}$
The $\left(\sum\limits_{i=1}^{m}[y^{(i)}=j]\right)$ term is constant, so we drop it as a proportionality constant (say, $C$). Therefore,
$\nabla^{2}_{W}E$ $\leq C\frac{k-1}{km}\nabla_{W}a_{j}^{[L-1]}$
$=C\frac{k-1}{km}\dfrac{\partial a_{j}^{[L-1]}}{\partial E}\dfrac{\partial E}{\partial w_{ij}}$
$\leq C\frac{(k-1)^{2}a_{j}^{[L-1]}}{k^{2}m^{2}}\dfrac{\partial a_{j}^{[L-1]}}{\partial E}$ using chain rule on Lemma 1
$=C\frac{(k-1)^{2}a_{j}^{[L-1]}}{k^{2}m^{2}}\frac{1}{\dfrac{\partial E}{\partial a_{j}^{[L-1]}}}$
Using $z^{[l]}=W^{[l]T}a^{[l-1]}+b^{[l]}$ and Lemma 1 gives us:
$\dfrac{\partial E}{\partial a_{j}^{[L-1]}}$ $=\dfrac{\partial E}{\partial z_{j}^{[L]}}\dfrac{\partial z_{j}^{[L]}}{\partial a_{j}^{[L-1]}}\leq\frac{(k-1)}{km}\lVert W^{[L]}\rVert$
from which (9) follows. ∎
Note: It is typical to use a regularization term in addition to the standard loss function above. In the supplementary material, we show that the addition of the regularizer only adds a constant term to the smoothness, and as such, it can be disregarded.
To report all this maths in another way, all that is required is to compute the norm of the weight matrix at the last layer. Of the $N_{1}$ configurations trained for one epoch, we choose the $N_{2}$ with the highest norms. This is because $k$ and $m$ are constant for a dataset, and so $\frac{k-1}{km}$ is simply a scaling factor.
4.3 Smoothness for Logistic Regression
Corollary 1.1 (Smoothness of logistic regression).
$\beta=\frac{k-1}{km}\sup\frac{\lVert X\rVert}{\lVert W\rVert}$
Proof.
Logistic regression can be viewed as a feedforward network with no hidden layers. Hence, the smoothness for logistic regression follows from Equation (9). ∎
To say that another way, the smoothness computation for logistic regression is particularly easy: after training the learner, the practitioner simply computes the norm of the input data and the weights after that single epoch of training and uses them to compute the smoothness.
4.4 Smoothness for Gaussian Naive Bayes
For the proof of the following, please see the supplementary material.
Theorem 2 (Smoothness for Gaussian Naive Bayes).
For Naive Bayes, the smoothness is given by
$\beta=\sup\lVert\boldsymbol{\Sigma}^{-1}\mathcal{A}G-\frac{1}{2}\boldsymbol{\Sigma}^{-1}\mathcal{A}\boldsymbol{\Sigma}^{-1}+G\mathcal{A}\boldsymbol{\Sigma}^{-1}\rVert$
where $\boldsymbol{\Sigma}$ is the covariance matrix, $\mathcal{A}$ is the fourth-order identity tensor, and $G=\boldsymbol{\Sigma}^{-1}\left((\boldsymbol{x}-\boldsymbol{\mu})(\boldsymbol{x}-\boldsymbol{\mu})^{T}+\frac{1}{2}\boldsymbol{\Sigma}\right)\boldsymbol{\Sigma}^{-1}$.
Notice that the smoothness is dependent only on the data $X$, means of the classes $\mu_{i}$, and the covariance matrix $\Sigma$), the latter two of which can also be derived from the dataset.
To say that another way, in practice, one would
  • Extract the means and covariance matrix from the dataset
  • Compute $G$, which simply requires us to compute the inverse of the covariance matrix and the deviations of each data point from the mean for the corresponding class
  • Use the intermediaries derived above to compute the smoothness.
It is worth noting that this is only true for Gaussian Naive Bayes. This is because our proof relies on the fact that Gaussian Naive Bayes is a special case of Gaussian Discriminant Analysis; this is not true for multinomial Naive Bayes.
4.5 Smoothness for other learners
While it is possible to use the methodology used in this paper to derive the smoothness for other classical learners, we leave that to future work. Right now we are experimenting with a non-parametric form of all the above where some recursive cluster method, or a decision tree learner, divides the data into a tree with small leaves. Then we report smoothness by computing the difference between neighbouring leaves and propagating that change up the tree using a center difference approximation:
$(\nabla^{2}E)(x)\approx\frac{E(x+h)-2E(x)+E(x-h)}{h^{2}}$
where $E$ is an impurity measure and the approximation error is $\mathcal{O}(h^{2})$.
Preliminary results are promising but, at this time, we have nothing definitive to report.
Figure 5: Summary of hyper-parameter space explored in this paper.
TABLE III: Dataset sizes for each of the tasks in this study. Numbers on brackets denote number of SE papers we could find that have explore these options (in a Google search of "hyper-parameter optimization" and "software engineering" 2014-2024).
Defect prediction [70]
Dataset #Train (%defects) #Test (%defects) #Features
camel 1819 (21) 442 (64) 20
ivy 352 (22) 352 (11)
log4j 244 (29) 205 (92)
synapse 379 (20) 256 (34)
velocity 410 (70) 229 (34)
xalan 2411 (38) 909 (99)
Static code warnings [72]
Dataset #Train #Test #Features
ant 28 7 329
cassandra 13 4 265
commons 17 5 127
derby 366 92 292
jmeter 14 4 287
lucene-solr 24 6 278
tomcat 147 37 284
Issue lifetime prediction [71, 43]
Dataset #Train #Test #Features
Eclipse 44,545 25,459 12
Firefox 44,800 25,201
Chromium 44,801 25,200
Bayesmark datasets
Dataset #Train #Test #Features
breast 455 114 30
digits 1438 359 64
iris 120 30 4
wine 142 36 13
diabetes 354 88 10
5 Experimental Rig
To test SMOOTHIE, we apply the following rig.
5.1 Data
Table III describes the datasets used in this study. Clearly, our study encompasses a variety of data shapes: both small and large, and both wide and narrow. In particular, for static code warnings, the number of features varies by dataset because of the data collection process from FindBugs; see [38] for a detailed report of the data cleaning process.
5.2 Learners
For our learners, we use the three for which we can currently calculate the smoothness of their loss function; i.e. feedforward networks, Naive Bayes, and logistic regression.
Bayesmark [59] is a benchmark suite used at the NeurIPS 2020 Black-Box Optimization Challenge. The datasets used in the challenge are available via OpenML [^6] https://openml.org, and are a part of the HPOBench [18] benchmark. Bayesmark, as a benchmarking tool, uses a normalized scoring system that discourages HPO methods that perform close to random search.
5.3 Hyper-Parameter Options
Our hyper-parameter options come from Tables III, and are summarized in Figure 5. We use these options since, as detailed in Table II, all those options have found favor with prior researchers. In Figure 5, branches with a solid line represent an AND, while dashed lines represent an OR. The learner-independent space has 60 options, and the learners can be configured in 100 ways. This yields 60 * 100 = 6,$0$ options in total. For the experiments on the Bayesmark datasets, we use the hyper-parameter space used by Bayesmark, which for feedforward networks optimized by Adam, has a size of 134M options [^7] https://github.com/uber/bayesmark/blob/master/bayesmark/sklearn_funcs.py#L91-L101.
Fortunately, many of these options have similar effects. As show below that within a random selection of 30, it is possible to find settings that significantly improve the performance of an off-the-shelf learner.
5.4 Hyper-Parameter Optimizers
For all our datasets, we compare against BOHB, DEHB, random search, and the current state-of-the-art for those tasks. For SMOOTHIE, we try all the learners whose smoothness value we derived; we use the best-performing learner as the base for random search, BOHB, and DEHB (described above in §3). In those trials, we compared to BOHB after $N_{1}$ and $N_{2}$ evaluations, and DEHB after $N_{1}$ evaluations. This gives both methods the fairest chance.
We briefly mention that although DEHB can be run in a distributed fashion and reduce runtime that way, we instead compare with it in a simpler, single-machine setup. This is since SMOOTHIE is also trivially parallelizable (since all trials in a stage are independent). We argue that this approach maintains fairness when comparing runtimes for all methods.
5.4.1 Train and Test Sets
For all experiments, we split the data into train and test partitions, and report results on the test set. We use the splitting policies seen in prior work:
  • For defect prediction, it is standard practice to train on data from versions $11\ldots n-$ 1 and test on version $n$ [2];
  • For issue lifetime prediction, we split the data 75/25;
  • For static code analysis, we split the data 80/20.
  • Bayesmark splits the dataset into 80/20 internally.
5.5 Evaluation Metrics
For each kind of data set, we use evaluation measures seen in prior papers that explored that data:
  • For issue lifetime prediction, as done in prior work [43] the class variable was divided into $k\in\{2,3\}$ equal frequency bins. We use the same metrics as [43], i.e. accuracy. If the true negative, false negative, false positive and true positive counts are $a,b,c,d$ then
[…]
  • For defect prediction, as done in prior work [63, 70], we used the F1 evaluation metric that commented not just on how many defects were were found but also, of the predicted defects, how many were true defects:
    $\begin{array}[]{rcccl}\mathrm{recall}&=&r&=d/(b+d)\\ \mathrm{precision}&=&p&=d/(c+d)\\ \mathrm{harmonic\;mean}&=&\mathit{F1}&=2rp/(r+p)\end{array}$
  • For static code false alarms, as done in prior work [72], we use recall, precision, and false alarm:
    $\mathrm{false\ alarm}=c/(a+c)$
To test the generality of SMOOTHIE of a broader range of data sets, we also used the Bayesmark dataset, which was previously used at NeurIPS 2020 in their Black-Box Optimization Challenge. The Bayesmark score scales the model performance based on how well random search does, after normalizing the mean over multiple repeats. The normalized mean score first calculates the performance gap between observations and the global optimum and divides it by the gap between random search and the optimum.
Dataset Method (Learner) #evals pd pf precision
ant Baseline (SVM) 0 0.25 0
SMOOTHIE (FF) 10 0.65 0 0.8
SMOOTHIE (NB) 10 0.6 0 0.8
SMOOTHIE (LR) 10 1 0 1
Random 10 0.2 0.5 0.5
BOHB (FF) 10 1 0.2 0.5
BOHB (FF) 50 0.6 1 0.6
DEHB (FF) 50 1 0.8 0.3
cassandra Baseline (SVM) 0.33 1 0.5
SMOOTHIE (FF) 10 0.83 0 0.75
SMOOTHIE (NB) 10 1 0.5 0.83
SMOOTHIE (LR) 10 1 0.67 0.33
Random 10 0 0.66 0
BOHB (FF) 10 1 0.67 0.25
BOHB (FF) 50 0 0 0
DEHB (FF) 50 1 0.67 0.75
commons Baseline (SVM) 1 1 0.6
SMOOTHIE (FF) 10 1 0.25 0.5
SMOOTHIE (NB) 10 0 0 0
SMOOTHIE (LR) 10 1 1 0.8
Random 10 0 0 0
BOHB (FF) 10 0.88 1 0.65
BOHB (FF) 50 1 1 0.2
DEHB (FF) 50 1 0.5 0.33
derby Baseline (SVM) 0.61 0.67 0.39
SMOOTHIE (FF) 10 0.74 0.46 0.44
SMOOTHIE (NB) 10 0.90 0.60 0.4
SMOOTHIE (LR) 10 0.59 0.06 0.8
Random 10 1.0 1.0 0.68
BOHB (FF) 10 0.93 0.84 0.68
BOHB (FF) 50 1 1 0.68
DEHB (FF) 50 0.9 0.68 0.38
jmeter Baseline (SVM) 0 0 0
SMOOTHIE (FF) 10 1 0 1
SMOOTHIE (NB) 10 1 0.33 0.5
SMOOTHIE (LR) 10 0.67 0 0.63
Random 10 0 1 0
BOHB (FF) 10 0 0 0
BOHB (FF) 50 0 0 0
DEHB (FF) 50 1 0.33 0.5
lucene-solr Baseline (SVM) 0 1 0
SMOOTHIE (FF) 10 0.5 0.25 0.5
SMOOTHIE (NB) 10 1 0.5 0.5
SMOOTHIE (LR) 10 1 0.25 0.67
Random 10 1 0.25 0.67
BOHB (FF) 10 0.75 0.5 0.33
BOHB (FF) 50 0 0 0
DEHB (FF) 50 1 0.5 0.5
tomcat Baseline (SVM) 0.73 0.77 0.39
SMOOTHIE (FF) 10 0.14 0.09 0.47
SMOOTHIE (NB) 10 0.43 0.13 0.63
SMOOTHIE (LR) 10 0.71 0.3 0.61
Random 10 0.43 0.5 0.59
BOHB (FF) 10 0.7 0.48 0.6
BOHB (FF) 50 1 1 0.62
DEHB (FF) 50 1 1 0.38
TABLE IV: RQ1 results: Static code warnings results. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6). For Algorithm 1, we use $N_{1}=50,N_{2}=10$. FF = Feedforward network. NB = Naive Bayes. LR = Logistic Regression.
5.6 Statistical Analysis
To compare multiple groups of data, we use a combination of the Kruskal-Wallis test (as used in recent SE literature [26, 53]) and the Mann-Whitney U-test [50] at a 5% significance level. In particular, we first use the Kruskal-Wallis H-test to determine if there is a statistically significant difference between the multiple groups. If there is no significant difference, we report all treatments as indistinguishable; if there is, we run pairwise Mann-Whitney U-tests to find treatments that are not statistically significantly different from the highest-scoring treatment (using a significance level of 0.05). We apply the Benjamini-Hochberg correction to adjust p-values (as endorsed by Farcomeni [20]) of the Mann-Whitney U-tests.
6 Results
Dataset Learner HPO method #evals Accuracy
eclipse (2) Baseline 61
Feedforward SMOOTHIE 5 69.3
Naive Bayes SMOOTHIE 5 63
Logistic SMOOTHIE 5 62.5
Best (Feedforward) Random 5 80.2
Best (Feedforward) BOHB 5 45.8
Best (Feedforward) BOHB 30 45.8
Best (Feedforward) DEHB 30 68.7
firefox (2) Baseline 67
Feedforward SMOOTHIE 5 69
Naive Bayes SMOOTHIE 5 66.6
Logistic SMOOTHIE 5 66.4
Best (Feedforward) Random 5 76.5
Best (Feedforward) BOHB 5 36
Best (Feedforward) BOHB 30 35.9
Best (Feedforward) DEHB 30 72.4
chromium (2) Baseline 63
Feedforward SMOOTHIE 5 62.2
Naive Bayes SMOOTHIE 5 63
Logistic SMOOTHIE 5 63.1
Best (Logistic) Random 5 67.1
Best (Logistic) BOHB 5 44.5
Best (Logistic) BOHB 30 44.4
Best (Logistic) DEHB 30 66.8
eclipse (3) Baseline 44
Feedforward SMOOTHIE 5 71.2
Naive Bayes SMOOTHIE 5 51.6
Logistic SMOOTHIE 5 51.2
Best (Feedforward) Random 5 36.3
Best (Feedforward) BOHB 5 29.2
Best (Feedforward) BOHB 30 46.2
Best (Feedforward) DEHB 30 71.2
firefox (3) Baseline 44
Feedforward SMOOTHIE 5 69.2
Naive Bayes SMOOTHIE 5 49.5
Logistic SMOOTHIE 5 49.3
Best (Feedforward) Random 5 46.1
Best (Feedforward) BOHB 5 32.9
Best (Feedforward) BOHB 30 44.4
Best (Feedforward) DEHB 30 70.5
chromium (3) Baseline 43
Feedforward SMOOTHIE 5 71.9
Naive Bayes SMOOTHIE 5 49.1
Logistic SMOOTHIE 5 49.8
Best (Feedforward) Random 5 53.5
Best (Feedforward) BOHB 5 27.5
Best (Feedforward) BOHB 30 44.6
Best (Feedforward) DEHB 30 72.9
TABLE V: RQ1 results: Issue lifetime prediction results. For Algorithm 1, we use $N_{1},N_{2}=30,5$. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6). Our baseline is DeepTriage [49]. In the dataset column, the number of classes the data is binned into is in parentheses.
Dataset Learner HPO method #evals F1
camel Baseline (GHOST) 90.9
Feedforward SMOOTHIE 5 90.9
Naive Bayes SMOOTHIE 5 90.9
Logistic SMOOTHIE 5 90.9
Best (NB) Random 5 90.9
Best (NB) BOHB 5 90.9
Best (NB) BOHB 30 90.9
Best DEHB 30 42.5
ivy Baseline (GHOST) 83.4
Feedforward SMOOTHIE 5 94
Naive Bayes SMOOTHIE 5 94
Logistic SMOOTHIE 5 94
Best (NB) Random 5 94
Best (NB) BOHB 5 94
Best (NB) BOHB 30 94
Best DEHB 30 31.2
log4j Baseline (GHOST) 65.9
Feedforward SMOOTHIE 5 79.6
Naive Bayes SMOOTHIE 5 79.6
Logistic SMOOTHIE 5 79.6
Best (NB) Random 5 79.6
Best (NB) BOHB 5 79.6
Best (NB) BOHB 30 79.6
Best DEHB 30 12.7
synapse Baseline (GHOST) 84.6
Feedforward SMOOTHIE 5 84.4
Naive Bayes SMOOTHIE 5 84.4
Logistic SMOOTHIE 5 84.4
Best (NB) Random 5 84.4
Best (NB) BOHB 5 74.5
Best (NB) BOHB 30 0
Best DEHB 30 56.2
velocity Baseline (GHOST) 50.3
Feedforward SMOOTHIE 5 55.4
Naive Bayes SMOOTHIE 5 79.8
Logistic SMOOTHIE 5 79.8
Best (NB) Random 5 79.8
Best (NB) BOHB 5 79.8
Best (NB) BOHB 30 79.8
Best DEHB 30 54.7
xalan Baseline (GHOST) 66
Feedforward SMOOTHIE 5 68.3
Naive Bayes SMOOTHIE 5 68.3
Logistic SMOOTHIE 5 68.3
Best (Logistic) Random 5 68.2
Best (Logistic) BOHB 5 0
Best (Logistic) BOHB 30 68.3
Best DEHB 30 45.5
TABLE VI: RQ1 results: Defect prediction results. For Algorithm 1, we use $N_{1},N_{2}$ =30, 5. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6).
TABLE VII: RQ1 results: Results on the NeurIPS 2020 Black-Box Optimization Challenge public datasets. For Algorithm 1, we use $N_{1},N_{2}=30,5$. Anything that ties with the statistically best results are shown in bold (and such ties are determined using the statistics of §5.6).
Dataset HPO method Score
breast Baseline (HEBO) 90.84
SMOOTHIE 94.32
Random search 90.45
BOHB 92.37
DEHB 91.72
digits Baseline (HEBO) 95.24
SMOOTHIE 89.51
Random search 87.39
BOHB 91.50
DEHB 96.87
iris Baseline (HEBO) 78.27
SMOOTHIE 87.64
Random search 82.98
BOHB 92.30
DEHB 80.78
wine Baseline (HEBO) 74.93
SMOOTHIE 93.34
Random search 76.59
BOHB 81.68
DEHB 81.73
diabetes Baseline (HEBO) 98.67
SMOOTHIE 98.65
Random search 97.16
BOHB 97.86
DEHB 99.40
We ran the experiments for defect prediction and issue lifetime prediction on a university Slurm cluster. For static code analysis, we used a Google Cloud N2 machine with 4vCPUs and 23GB RAM. For the Bayesmark datasets, we ran the experiments on a Google Cloud C3 VM.
Using those results, we explore three research questions:
  • RQ1: What are the benefits of smoothing (measured in metrics like F1, recall, etc)?
  • RQ2: What are the costs of smoothing (measured in CPU)?
  • RQ3: Should we use different kinds of algorithms for our SE data? Note that this third question addresses the issues raised in our introduction.
6.1 RQ1: What are the benefits of smoothing?
Table IV shows our results on static code analysis false alarms. Recall that the task here was to recognize static code warnings that, in the historical log, were ignored by developers. In the table, gray denotes “best results in this column, for that dataset“. Note that SMOOTHIE (and more specifically, SMOOTHIE (FF) has much more gray than anything else. More specifically: all the non-SMOOTHIE methods have far more white cells (i.e. non best) than the SMOOTHIE methods. (Table VIII shows the exact win/tie counts).
Table V shows our results on issue lifetime prediction. Recall that the task here is to classify a GitHub issue according to how long it will take to resolve. As done in prior work [43], for 3-class classification, the number of days taken to fix bugs is binned into 0 to 1 days, 2 to 5 days, and 5 to 100 days. For 2-class, the number of days taken to fix bugs is binned into less than (or greater than) 4 days for Firefox, less than (or greater than) 5 days for Chromium, and less than (or greater than) 6 days for Eclipse. Note that:
  • For two class prediction (eclipse(2), firefox(2), chromium(2)) SMOOTHIE does not perform well (as shown in Table VIII, 1 tie and 2 losses). These datasets have the highest smoothness (i.e., sharpest minima) over all our SE datasets: the 2-class Chromium, Eclipse, and Firefox datasets had smoothness values of 9.4, 14.9, and 13.8 respectively. As discussed in RQ3 (Figure 7), we expect SMOOTHIE to perform poorly on datasets with such a high smoothness. This explains why SMOOTHIE did not work well for the 2-class binned version of the data.
  • For three-class problems (eclipse(3), firefox(3), chromium(3)) , SMOOTHIE wins in all cases.
TABLE VIII: RQ1 results (summary): from 20 repeats over 24 data sets.
Frequency counts of:
Datasets wins ties losses
Static code warnings 4 3 0
Issue lifetime prediction (binary) 0 1 2
Issue lifetime prediction (multi-class) 0 3 0
Defect prediction 0 6 0
Bayesmark 2 1 2
Table VI shows defect prediction results. Recall that this task imports static code features and outputs a prediction of where bugs might be found. It turns out that defect prediction is our simplest task (evidence: all methods achieve the same best ranked results); defect prediction may not be a useful domain to compare optimizers.
Table VII shows our results for the NeurIPS 2020 Black- Box Optimization Challenge public dataset. We include these results here just to study how well SMOOTHIE performs on non-SE data.
Table VIII summarizes the RQ1 results from Table IV to Table VII. In that table, two methods tie if, they achieve the same statistical rank (using the methods of §5.6). Otherwise, a win is when one measure has a different rank to another and one measure has better median values (where “better” means “less“ for false alarms and “greater” for other metrics).
In summary, from Table VIII, we say:
[…]
Figure 6: RQ2 Results: Runtime results on the chromium-2 class dataset.
6.2 RQ2: What are the costs of smoothing?
Figure 6 shows the runtimes for our method. We take care to distinguish the overheads associated with finding smoothness, versus the time required for another process to use smoothness to make a recommendation about what are good hyperparameter options. All algorithms were measured on a Google Cloud n1-standard-8 VM with a P4 GPU.
As shown in Figure 6, the time required to compute Get-Smoothness smoothness is negligible. Those results are then used by SMOOTHIE for its decision making (and here, SMOOTHIE repeats its processing $N_{1}+N_{2}=30+5$ times). Even with those repeats, SMOOTHIE is 4.7 times faster than BOHB and 8.3 times faster than DEHB.
Figure 6 suggests that further optimizations might be possible if SMOOTHIE could reduce how many repeats its executes. This is a matter for further work, in another paper.
[…]
6.3 RQ3: Should we use different kinds of algorithms for our SE data?
Apart from performance and runtime details, what does the above mean for the challenge problem sketched out in the introduction to this paper? To answer that question, we first recall Figure 3 where the mean smoothness for our SE and AI data has values 2 and 5, respectively.
Turning now to Figure 7, that figure reports the probability of SMOOTHIE outperforming BOHB, measured as the fraction of datasets for which SMOOTHIE outperforms BOHB (measured in terms of the F1 metric) [^8] More precisely, $\forall i>0,f(x_{i})=\frac{\mathbbm{1}\{(x_{i-1}\leq x\leq x_{i})\land(\text{SMOOTHIE}>\text{BOHB})\}}{\mathbbm{1}\{(x_{i-1}\leq x\leq x_{i})\}}$
We note that, in those results, different algorithms are the best choice for low and high smoothness values:
  • SMOOTHIE usually wins at the low values typically associated with SE data;
  • While BOHB, DEHB usually wins at the higher values, seen only in the non-SE data.
[…]
6.4 Sensitivity Analysis
We also conduct a sensitivity analysis over the parameters $N_{1}$ and $N_{2}$ on the static code analysis datasets (excluding maven, due to its smaller size). Figure 8 shows these results.
On the x-axis is the number of initial evaluations ($N_{1}$ in this paper) performed, and each line corresponds to one choice of full runs performed ($N_{2}$ in this paper). We start with a minimum of $N_{1}=20$, and go up to $N_{1}=75$, and test with $N_{2}\in\{5,10,15\}$. The y-axis shows the mean normalized regret over evaluations, where lower is better. This normalized regret is calculated as
Normalized $\text{Normalized regret}=1-\frac{p_{t}}{p_{20}}$
where $p_{t}$ is the performance (measured as the difference between recall and false positive rate). We note that as $N_{1}$ increases, the normalized regret increases. This is because $N_{2}$ stays fixed for each line, so the highest smoothness changes as $N_{1}$ increases. Further, we note that up to $N_{1}\approx 45$, the normalized regret for $N_{2}=5$ is low (and lower than $N_{2}=15$). Past this point, the normalized regret is higher. However, it is worth noting that moving from $N_{2}=5$ to $N_{2}=10,15$ incurs a 2-3x runtime penalty.
This figure shows that $N_{1}=30,N_{2}=5$ are reasonable default values for these parameters.
7 Discussion
7.1 Why Does SMOOTHIE Work?
Figure 7 raises three questions:
  • Why does BOHB fail for small $\beta$-smoothness values (i.e. the low $x$ values of Figure 7)?
  • Why do those same flatter regions help SMOOTHIE?
  • Why does SMOOTHIE fail for for very bumpy regions (i.e. the high $x$ values of Figure 7)?
Figure 7: Probability of SMOOTHIE outperforming (or matching) BOHB, DEHB, and random as a function of smoothness
Figure 8: Sensitivity analysis for the static code analysis datasets, excluding maven.
We offer the hypothesis that algorithms like BOHB need a bumpy landscape to distinguish between different parts of the data, in order to decide what regions to explore, so BOHB prefers a bumpier space.
But if BOHB fails in flat regions, why would SMOOTHIE succeed in the same flatter regions? It is well-established that learning algorithms (such as SGD or Adam) have a trade-off between training rates and generalization [21, 34, 45, 16, 39, 65, 47, 74]. Steep landscape decrease training rates since at any point there is an obvious better place to jump to. But flatness helps generalization (i.e. scores on hold-out test data). Many theorists [40, 34, 33, 17, 37, 66, 35, 54, 32, 12, 56] comment that the flatter the loss landscape, the more generalizable the conclusion (since for any given conclusion, little changes in adjacent regions). To see this, consider the what happens when the loss landscape is not flat: what is learned in one region might change significantly in the adjacent region.
In short, we believe that BOHB does not explicitly bias for generalization, only for test performance, and SMOOTHIE gets higher scores on test data by better exploiting generalization over flatter terrains.
As to why SMOOTHIE fails for bumpy spaces, we say its design trade-offs do not favor such a space, for the following reason. We conjecture that SMOOTHIE flattens the loss function by some ratio of, say, 50%. If this were true then when $\beta$-smoothness has a large value of, say, 5.6 (see Figure 9(d).d), SMOOTHIE can only flatten the landscape to .5*5.6=2.8 (see Figure 9(d).c, which is still quite steep). On the other hand, if we start near the mean smoothness of our SE data (see Figure 9(d).b) then a 50% flattening takes us into a quite flat region (see Figure 9(d).a) where models generalize better.
(a) $z=0.3(x^{2}+y^{2}),\beta=0.85$
(b) $z=0.6(x^{2}+y^{2}),\beta=1.7$
(c) $z=x^{2}+y^{2},\beta=2.8$
(d) $z=2(x^{2}+y^{2}),\beta=5.6$
Figure 9: An illustration of sharp minima forming with higher smoothness. Figure 9(d)d is the upper limit of the sharpness that SMOOTHIE can handle (see Figure 7).
All figures generated using Wolfram Alpha.
7.2 Threats to Validity
In this section, we discuss various biases that might affect the final results, and how our experimental setup attempts to mitigate them.
Sampling bias. Sampling bias is an issue for experimental papers that perform evaluation on a limited set of data points. We believe this is mitigated by testing our method on 24 datasets against a state-of-the-art optimizer. Our code is open-source, and we invite other researchers to extend our results to more datasets.
Evaluation bias. Evaluation bias threatens papers that use a limited number of metrics to evaluate performance. It is important to ensure all comparisons are fair and provide a broad view of the performance. By using the same metrics as prior work to compare the methods in this paper, we demonstrate that our method can achieve state-of-the-art results across a variety of tasks and metrics.
Learner bias. Another source of bias is the learners used in the study. Currently, there are many algorithms for classification (which this paper focuses on). We developed our theory for three learners and tested them across all our datasets in an effort to mitigate this bias. In future work, we will implement the finite difference approximation (cf. Section 4.5) which open up the theory to arbitrary algorithms.
Comparison bias. There are many approaches for the tasks we studied in this paper, and it would be difficult to compare to all of them. As such, we decided to compare our work to the current state-of-the-art, as well as a state-of-the-art hyper-parameter optimization algorithm. This ensures that our results are compared against the best among the algorithms in the literature.
Order bias. Order bias has to do with the order of samples affecting performance. For defect prediction, we ensured that the test set is from a newer version than the training set (since that data comes with time stamps). For all other datasets, we mitigate this bias by randomly splitting the dataset into train and test sets, and repeating our results 20 times.
Statistical validity. Especially with stochastic algorithms, it is important to ensure that any performance differences cannot be explained by noise. We repeat all our experiments 20 times and use adjusted p-values from a widely-used statistical test to declare wins, ties, and losses.
8 Limitations and Future Work
The major limitation of this work is the choice of measurement of smoothness, and the use of smoothness as a heuristic. For example, having derived the Hessian, the condition number is an alternate measure of the “nice nature” of the landscape [11]. Another clear limitation is that this approach is limited to learners for which a smoothness can be derived. Especially in the case where the Hessian may not exist (e.g., if the loss function is not twice-differentiable such as the mean absolute error loss), this approach fails. Nevertheless, we have shown successful application in a classical learner that is not gradient-based, Naive Bayes. In that case, we used the negative log-likelihood as a surrogate for the loss function. We postulate that this should be possible for other learners too, such as the dual formulation in SVMs or the negative log-likelihood in decision trees. This issue would a frutiful direction for future work. Finally, we reiterate that SMOOTHIE performs poorly on datasets with high smoothness, as evidenced by our results on the 2-class issue close time prediction results.
Another avenue of future work would be replacing the random sampling at the beginning of the approach (Algorithm 1, Line 1) with a smarter selection approach. For example, Bayesian optimization could be used in a loop to search for configurations with high smoothness values. Alternatively, one could use multiple metrics for smoothness, such as the condition number of the Hessian, and use a multi-objective optimization algorithm.
Finally, this paper only explored three learners. In future work, we plan to extend this to more tasks and learners such as convolutional networks.
9 Conclusion
One of the black arts of software analytics is how to configure our data mining systems. Hyper-parameter optimization is often implemented as a blind search through a large space of options (like randomly hitting a broken car with a hammer, until the car starts working).
This paper offers a novel theory that can better guide the design hyper-parameter optimizers:
Hyper-parameter optimization for SE is best viewed as a “smoothing” function for the decision landscape since, after optimization, the loss function gets “smoother’ (i.e. on average, fewer changes in any local region).
This theory predicts that methods that maximize the smoothness should also improve the hyper-parameter optimization. This prediction was checked on on 24 datasets across a variety of tasks. As predicted, SMOOTHIE, achieves state-of-the-art performance and outperforms other popular HPO methods. One reason to endorse this theory is the time required to compute smoothness is negligible.
We show that a hyper-parameter optimizer called SMOOTHIE, which is based on this theory:
  • Works well for flatter landscapes;
  • But, compared to standard AI tools like BOHB, fails for much bumpier spaces.
  • On the other hand the standard AI tools fail for the flatter spaces.
This is a significant observation since, as we shown in Figure 3, SE data can be much flatter than the data usually processed by the AI tools.
Based on that results, we conclude that SE data can be different to other kinds of data; and those differences means that, sometimes, we should use different kinds of algorithms for our data. Hence it is unwise to always use these AI algorithms in their default “off-the-shelf” configuration.
Acknowledgements
This work was partially funded by an NSF grant #1908762.